Problem 5. Show that an element ū is a unit in R/PR, and find its inverse. (a) (15 pts) R=F3 [C], p= x4 + 2x +1, u = x2 +1. (Only B
Answer #1

Ipablems Given R=2GI, P=8771, U=1-21 consider 2010 AS 8471 E (8+7i) then 8+71 + (8771) = 0+ (8+70) at bi +18+71) atbi (say )

= 2 2 114 2 5x 23= 2X57 = 2 x 57 ». (5x23,57 Toe by eavation to 4 and D 224 D (úxit2ilx 25x57 at by and 5 X 2357*(1+2;) = ī 2